OpenStudy (anonymous):

Give an example of a matrix which is its own inverse (that is, where A^-1 = A)

OpenStudy (anonymous):

the identity matrices

OpenStudy (anonymous):

The answer is Many answers ex: -10 9 -11 10 But I don't understand... how did they get this answer?

OpenStudy (anonymous):

What about the identity matrices?

OpenStudy (anonymous):

the inverse of an identity is an identity itsself. and an identity times any matrix gives the same matrix. so any identity matrix is a solution

OpenStudy (anonymous):

no thats wrong

OpenStudy (anonymous):

\[\left[\begin{matrix}1 & 0 \\ 0 & 1\end{matrix}\right]\]

OpenStudy (anonymous):

I know the identity matrix is 1 0 0 1

OpenStudy (maheshmeghwal9):

oops sorry my bad a mistake in haste:/

OpenStudy (anonymous):

But I don't understand how to do this :(

OpenStudy (anonymous):

I understand what an identity matrix is, but I don't know how to come up with a matrix myself

OpenStudy (anonymous):

that is its own inverse

OpenStudy (anonymous):

I guess it's any matrix that has 1 as its determinant?

OpenStudy (anonymous):

-1

OpenStudy (maheshmeghwal9):

\[A=A^{-1}.\]\[\implies \frac{A}{A^{-1}}=1I.\]\[\implies A \times A^1=A \times A=A^2=1I.\]\[\color{red}{\implies A= \sqrt{1I}=1\sqrt{I}=\sqrt{I}=I.}\]Which is an identity matrix.